# Dept. of Informatics King’s College London London, UK

### v1

@@ 4.1.2 Descriptive (ABM) Models of Hybrid Societies for DAMAS
Regarding the related work in MAS and ABM on distributed deception as we have defined it here, the closest works would be the one on evolution of deception in hybrid societies by Sarkadi et al. (2021a); Sarkadi and Lewis (2024) who model large-scale and long-term interactions as public-goods-games, and the work on the profitability of incompetence by Staab and Caminada (2010) who define artificial agents that bullshit their way through society in order to maintain the view that they (the agents) are competent.
@@
One-Way Deception Sakama and Caminada (2010); Sakama et al. (2010); Caminada (2009); Sakama et al. (2015); De Rosis et al. (2003); Christian and Young (2004); Panisson et al. (2018a); Sarkadi et al. (2019b)
Counter-Deception Sarkadi et al. (2019a)*
Self-Deception Jones (2015), Chen (2021)*
Distributed Deception Staab and Caminada (2010); Sarkadi et al. (2021a); Sarkadi and Lewis (2024)
@@ 4.2 Explaining Deception - Hybrid Argumentation Tools for DAMAS
For the interaction and descriptive models of DAMAS to be used in real-world analysis and decision making - in a meaningful manner - DAMAS should provide analysts (end users, not just AI scientists and experts) with some means or methods to explain and interpret the ABM and MAS dynamics. Remember that in history and Intelligence Analysis, to understand deception in a meaningful manner, deception causation and causation prevention need to be understood as part of an overarching narrative Ferris (1989).
In deception analysis, counterfactual or hypothetical reasoning deals with questions that are relevant for establishing event causation and event causation prevention. A promising approach in AI to address these questions is to use argumentation-enabled tools, which could be integrated in DAMAS for automated reasoning. Cybersecurity research is already looking at how the area of argumentation can be used to perform this type of analysis for tasks such as cyber attribution Simari (2019). Argumentation-based approaches can also be used in criminal forensics as proposed by Bex and Verheij (2010) and backed by Zenker (2019). The implementation of such support tools for intelligence analysis is not new and is indeed feasible, as previous work by Toniolo et al. (2015) demonstrates.
Given these capabilities of argumentation-enabled tools to facilitate explanations, we recommend to include in DAMAS Bex’s hybrid approach to argumentation and storytelling to explain causation and causation prevention as a narrative such that it is intelligible for analysts (see Fig. 2). Bex (2011)’s approach has been designed to both represent causal chains of a main story and to use arguments to anchor the main story’s sub-stories in evidence, a process named anchoring which results in explanations. By applying this approach, one does not only find an arbitrary story for a causal chain of events, but is able to select the best main story (out of several viable ones) that is composed of several sub-stories, e.g., perform the inference to the best explanation (IBE), by explaining through arguments how the main story’s sub-stories are backed by evidence.
@@
Representing multiple levels of abstraction should help analysts take into account the complete and partial knowledge of the agents involved in the respective scenarios - interaction MAS level, as well as the effect of the social factors that influence deception parameters - descriptive ABM level. Subsequently, analysts could describe and contrast the cognitive and communicative differences, while comparing similarities, that follow from the representations and behaviour of deceptive and non-deceptive agents. For example, it should be possible for an analyst to distinguish a lie from a lie with deceptive intent Chisholm and Feehan (1977), if components of deception could be described by the DAMAS modules by taking into account relevant theories and taxonomies of deceptive communication. Additionally, it is crucial for the analysts to be able to say if there is any intentional, unintentional, or even self-deception happening Whaley (1982), and be able to identify whether or not ‘banal’ deception is the case - deception that arises from using a technology that does not deceive on its own, but it is designed to be deceptive in a certain context Natale et al. (2021). A good way to identify banal deception in hybrid societies would be to look at cases where deception is caused by semi-autonomous machines which cannot be attributed intention Masters et al. (2021b), and represent these in DAMAS.
@@
Using Bex’s approach, the observable behaviour of agents inside DAMAS could be used to test whether a potential viable main story (a complex hypothesis) about their behaviour counts as an explanation of their behaviour. How well the story is anchored in the evidence could also be observed in DAMAS, e.g., how many items of evidence confirm or falsify the story. Furthermore, the story could be used to confirm or falsify how certain events in the real world might have unfolded, but more importantly it could explain how they might have unfolded or not if something else were the case. These are precisely the sorts of explanations of agent behaviour that human minds try to figure out in socio-cognitive interactions Malle (2006). One must also be careful about grounding the representation in relevant theories, as selecting bias-prone theories for this purpose leads to errors in analysing deception using AI methods Sánchez-Monedero and Dencik (2020); Stark and Hutson (2021). Anyways, Bex’s hybrid approach could also be used to find biases and inconsistencies inside DAMAS w.r.t. the theories that have been used to implement the MAS models. Perhaps there is considerable evidence outside the MAS that indicates that the agents’ behaviour should have been different. DAMAS would allow then the identification of the underlying cause for the unexpected behaviour, at which level of abstraction it happens, and whether it could be calibrated inside the model that represents that particular level of abstraction. Subsequently, following the feedback of analysts, AI experts could directly revise the underlying AI model according to a new evidence-based theory and re-implement it inside the framework. This process would be facilitated by DAMAS’s scalability and modularity.
@@
Stories can change based on new evidence, new characters, or new rules, which must be considered by DAMAS. A story can represent how a new agent character enters the story, or how the chain of events moves from one location (domain) to another, or perhaps the narrative style of the story allows for referring back to evidence or arguments that have a certain temporal property (using information available at a different time), e.g. something happened in the past that affected the current chain of events, or perhaps if this chain of events continues in this direction, something relevant to the story will happen in the future. Again, this would be facilitated by following the principles of scalability and modularity of designing DAMAS models in tune with the open nature of MAS systems van Eijk et al. (1999); Pinyol and Sabater-Mir (2013) where new types of agents or software tools are added or removed from the system.
@@
According to Miller (2019), explanations do not consist solely of using abductive reasoning to find causal relations, but they must also include a social process, which refers to the way of convening the knowledge to be exchanged between explainer and explainee. Regarding MAS tools, socially-aware AI meta-agents can be modelled and engineered as part of DAMAS so that they can automatically narrate various chains of complex events to analysts. DAMAS models that enhance abductive reasoning should be designed as to enable these socially-aware AI meta-agents to extract narratives directly from the agent dynamics of the MAS. Adding argumentation capabilities to socially-aware meta-agents could help analysts also engage in critical and hypothetical thinking about possible narratives using question-answering games. It is here where research in dialogue argumentation games for agent social interaction can provide mechanisms for reaching sound conclusions McBurney and Parsons (2009). The cooperation between humans and artificial agents would result in a better understanding of deceptive scenarios through dialectical reasoning. For instance, the approaches described in Shvo et al. (2020); Panisson et al. (2018b); Sarkadi et al. (2018); Mosca and Such (2021); Mosca et al. (2020) could be combined to model the social process between artificial storytelling meta-agents and their interlocutors (explainees). These approaches seem appropriate, because most humans prefer to explain agent behaviour in a practical reasoning manner by referring to beliefs, desires, intentions etc. Malle (1999, 2006). Storytelling meta-agents could model the minds of the explainees in order to narrate complex chains of events in an efficient manner - achieving the goal of explaining the phenomenon itself by taking into account the social context in which they explain the phenomenon. Conclusively, by interacting with a usable DAMAS, the analysts’ workload and tedious tasks, such as manually performing SATs, would be reduced and simplified through the interactions with socially-explainable AI agents.
@@ 5 Conclusion
Next, we argued that if we aim to model deception with AI techniques, then we must adopt a holistic socio-cognitive perspective because otherwise we would fall into the trap of scientific reductionism. We further argued that reductionism in AI is not in itself bad, but that in order to build a holistic understanding, then the existing AI approaches must be integrated as part of an overarching MAS framework. Our argument echoes what has been proposed about 20 years ago regarding the development of a theory to understand the role of trust and deception in virtual societies, notably by Castelfranchi (2000); Castelfranchi and Tan (2001, 2002). The reason we have reinforced this perspective is because despite the significant number of contributions on socio-cognitive modelling of trust, the socio-cognitive modelling of deception has not had the same momentum. Having surveyed the literature on deception modelling in AI, we believe it is now the time for the next stage in deception modelling and analysis.
@@
Finally, we discussed the impact of being able to understand deception holistically in hybrid societies and what are the ethical considerations of modelling deceptive AI. We pointed out the risk of reaching a Tragedy of the Digital Commons Greco and Floridi (2004) and discussed the need of developing accountable, responsible and transparent deceptive AI w.r.t. ethical guidelines, such as the guidelines of HLEG for trustworthy AI HLEG, in AI (2019) and the ones of IEEE on Ethically Aligned Design Chatila and Havens (2019). We then presented two arguments that defend the modelling and development of trustworthy deceptive AI for the benefit of society under the condition of following ethical principles of such as value-alignment, transparency, and scientific discovery.

### v2

@@ 4.1.2 Descriptive (ABM) Models of Hybrid Societies for DAMAS
Regarding the related work in MAS and ABM on distributed deception as we have defined it here, the closest works would be the one on evolution of deception in hybrid societies by (Sarkadi et al., 2021a; Sarkadi and Lewis, 2024) who model large-scale and long-term interactions as public-goods-games, and the work on the profitability of incompetence by (Staab and Caminada, 2010) who define artificial agents that bullshit their way through society in order to maintain the view that they (the agents) are competent.
@@
One-Way Deception (Sakama and Caminada, 2010; Sakama et al., 2010; Caminada, 2009; Sakama et al., 2015; De Rosis et al., 2003; Christian and Young, 2004; Panisson et al., 2018a; Sarkadi et al., 2019b)
Counter-Deception (Sarkadi et al., 2019a)*
Self-Deception (Jones, 2015), (Chen, 2021)*
Distributed Deception (Staab and Caminada, 2010; Sarkadi et al., 2021a; Sarkadi and Lewis, 2024)
@@ 4.2 Explaining Deception - Hybrid Argumentation Tools for DAMAS
For the interaction and descriptive models of DAMAS to be used in real-world analysis and decision making - in a meaningful manner - DAMAS should provide analysts (end users, not just AI scientists and experts) with some means or methods to explain and interpret the ABM and MAS dynamics. Remember that in history and Intelligence Analysis, to understand deception in a meaningful manner, deception causation and causation prevention need to be understood as part of an overarching narrative (Ferris, 1989).
In deception analysis, counterfactual or hypothetical reasoning deals with questions that are relevant for establishing event causation and event causation prevention. A promising approach in AI to address these questions is to use argumentation-enabled tools, which could be integrated in DAMAS for automated reasoning. Cybersecurity research is already looking at how the area of argumentation can be used to perform this type of analysis for tasks such as cyber attribution (Simari, 2019). Argumentation-based approaches can also be used in criminal forensics as proposed by (Bex and Verheij, 2010) and backed by (Zenker, 2019). The implementation of such support tools for intelligence analysis is not new and is indeed feasible, as previous work by (Toniolo et al., 2015) demonstrates.
Given these capabilities of argumentation-enabled tools to facilitate explanations, we recommend to include in DAMAS Bex’s hybrid approach to argumentation and storytelling to explain causation and causation prevention as a narrative such that it is intelligible for analysts (see Fig. 2). (Bex, 2011)’s approach has been designed to both represent causal chains of a main story and to use arguments to anchor the main story’s sub-stories in evidence, a process named anchoring which results in explanations. By applying this approach, one does not only find an arbitrary story for a causal chain of events, but is able to select the best main story (out of several viable ones) that is composed of several sub-stories, e.g., perform the inference to the best explanation (IBE), by explaining through arguments how the main story’s sub-stories are backed by evidence.
@@
Representing multiple levels of abstraction should help analysts take into account the complete and partial knowledge of the agents involved in the respective scenarios - interaction MAS level, as well as the effect of the social factors that influence deception parameters - descriptive ABM level. Subsequently, analysts could describe and contrast the cognitive and communicative differences, while comparing similarities, that follow from the representations and behaviour of deceptive and non-deceptive agents. For example, it should be possible for an analyst to distinguish a lie from a lie with deceptive intent (Chisholm and Feehan, 1977), if components of deception could be described by the DAMAS modules by taking into account relevant theories and taxonomies of deceptive communication. Additionally, it is crucial for the analysts to be able to say if there is any intentional, unintentional, or even self-deception happening (Whaley, 1982), and be able to identify whether or not ‘banal’ deception is the case - deception that arises from using a technology that does not deceive on its own, but it is designed to be deceptive in a certain context (Natale et al., 2021). A good way to identify banal deception in hybrid societies would be to look at cases where deception is caused by semi-autonomous machines which cannot be attributed intention (Masters et al., 2021b), and represent these in DAMAS.
@@
Using Bex’s approach, the observable behaviour of agents inside DAMAS could be used to test whether a potential viable main story (a complex hypothesis) about their behaviour counts as an explanation of their behaviour. How well the story is anchored in the evidence could also be observed in DAMAS, e.g., how many items of evidence confirm or falsify the story. Furthermore, the story could be used to confirm or falsify how certain events in the real world might have unfolded, but more importantly it could explain how they might have unfolded or not if something else were the case. These are precisely the sorts of explanations of agent behaviour that human minds try to figure out in socio-cognitive interactions (Malle, 2006). One must also be careful about grounding the representation in relevant theories, as selecting bias-prone theories for this purpose leads to errors in analysing deception using AI methods (Sánchez-Monedero and Dencik, 2020; Stark and Hutson, 2021). Anyways, Bex’s hybrid approach could also be used to find biases and inconsistencies inside DAMAS w.r.t. the theories that have been used to implement the MAS models. Perhaps there is considerable evidence outside the MAS that indicates that the agents’ behaviour should have been different. DAMAS would allow then the identification of the underlying cause for the unexpected behaviour, at which level of abstraction it happens, and whether it could be calibrated inside the model that represents that particular level of abstraction. Subsequently, following the feedback of analysts, AI experts could directly revise the underlying AI model according to a new evidence-based theory and re-implement it inside the framework. This process would be facilitated by DAMAS’s scalability and modularity.
@@
Stories can change based on new evidence, new characters, or new rules, which must be considered by DAMAS. A story can represent how a new agent character enters the story, or how the chain of events moves from one location (domain) to another, or perhaps the narrative style of the story allows for referring back to evidence or arguments that have a certain temporal property (using information available at a different time), e.g. something happened in the past that affected the current chain of events, or perhaps if this chain of events continues in this direction, something relevant to the story will happen in the future. Again, this would be facilitated by following the principles of scalability and modularity of designing DAMAS models in tune with the open nature of MAS systems (van Eijk et al., 1999; Pinyol and Sabater-Mir, 2013) where new types of agents or software tools are added or removed from the system.
@@
According to (Miller, 2019), explanations do not consist solely of using abductive reasoning to find causal relations, but they must also include a social process, which refers to the way of convening the knowledge to be exchanged between explainer and explainee. Regarding MAS tools, socially-aware AI meta-agents can be modelled and engineered as part of DAMAS so that they can automatically narrate various chains of complex events to analysts. DAMAS models that enhance abductive reasoning should be designed as to enable these socially-aware AI meta-agents to extract narratives directly from the agent dynamics of the MAS. Adding argumentation capabilities to socially-aware meta-agents could help analysts also engage in critical and hypothetical thinking about possible narratives using question-answering games. It is here where research in dialogue argumentation games for agent social interaction can provide mechanisms for reaching sound conclusions (McBurney and Parsons, 2009). The cooperation between humans and artificial agents would result in a better understanding of deceptive scenarios through dialectical reasoning. For instance, the approaches described in (Shvo et al., 2020; Panisson et al., 2018b; Sarkadi et al., 2018; Mosca and Such, 2021; Mosca et al., 2020) could be combined to model the social process between artificial storytelling meta-agents and their interlocutors (explainees). These approaches seem appropriate, because most humans prefer to explain agent behaviour in a practical reasoning manner by referring to beliefs, desires, intentions etc. (Malle, 1999, 2006). Storytelling meta-agents could model the minds of the explainees in order to narrate complex chains of events in an efficient manner - achieving the goal of explaining the phenomenon itself by taking into account the social context in which they explain the phenomenon. Conclusively, by interacting with a usable DAMAS, the analysts’ workload and tedious tasks, such as manually performing SATs, would be reduced and simplified through the interactions with socially-explainable AI agents.
@@ 5 Conclusion
Next, we argued that if we aim to model deception with AI techniques, then we must adopt a holistic socio-cognitive perspective because otherwise we would fall into the trap of scientific reductionism. We further argued that reductionism in AI is not in itself bad, but that in order to build a holistic understanding, then the existing AI approaches must be integrated as part of an overarching MAS framework. Our argument echoes what has been proposed about 20 years ago regarding the development of a theory to understand the role of trust and deception in virtual societies, notably by (Castelfranchi, 2000; Castelfranchi and Tan, 2001, 2002). The reason we have reinforced this perspective is because despite the significant number of contributions on socio-cognitive modelling of trust, the socio-cognitive modelling of deception has not had the same momentum. Having surveyed the literature on deception modelling in AI, we believe it is now the time for the next stage in deception modelling and analysis.
@@
Finally, we discussed the impact of being able to understand deception holistically in hybrid societies and what are the ethical considerations of modelling deceptive AI. We pointed out the risk of reaching a Tragedy of the Digital Commons (Greco and Floridi, 2004) and discussed the need of developing accountable, responsible and transparent deceptive AI w.r.t. ethical guidelines, such as the guidelines of HLEG for trustworthy AI (HLEG, in AI, 2019) and the ones of IEEE on Ethically Aligned Design (Chatila and Havens, 2019). We then presented two arguments that defend the modelling and development of trustworthy deceptive AI for the benefit of society under the condition of following ethical principles of such as value-alignment, transparency, and scientific discovery.